rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#5494 — FS#9415 — BHS : Test switch CDP

Attached to Project— Dedicated servers
Modernization
BHS1
CLOSED
100%
Tomorrow at 27/09/2013 Friday at 9:00 EST / 15.00 CET, we will perform a switch test at CDP1 and CDP2 to validate the modifications we made.

No cut-off is expected, UPS will take over the time of the changeover.
Date:  Saturday, 28 September 2013, 01:57AM
Reason for closing:  Done
Additional comments about closing:  Tests are done successfully.